solve please 7x+4=x+16
Answer by therei(2)

this is a linear equation..to solve for x..you just combine terms involving the variable (x) to the left hand side of the equation & all the constant to the right -hand side, that is .. 7x-x = 16- 4 ( make sure that if you transpose you change the sign of the term). Next simplify..
6x = 12, next divide both sides of the equation by the coefficient of x......6x/6 = 12/6
x = 2